OpenStudy (anonymous):

Simplify quantity 12 x minus 96 over quantity x squared minus 4x minus 32 all over quantity 6 x squared minus 6 x minus 12 over 3 x plus 3

OpenStudy (anonymous):

\[\frac{ \frac{ 12x-96 }{ x^2-4x-32 } }{ \frac{ 6x^2-6x-12 }{ 3x+3 } }\]

OpenStudy (anonymous):

\[\frac{ 12\cancel{(x-8)} }{ \cancel{(x-8)}(x+4) }\times \frac{ 3\cancel{(x+1)} }{ 6\cancel{(x+1)}(x-2) }\]

OpenStudy (anonymous):

\[\frac{ \cancel{12}6 }{ x+4 }\times \frac{ \cancel{3}1 }{ \cancel{3}\cancel{2}(x-2) }=\frac{ 6 }{ (x+4)(x-2) }\]

OpenStudy (anonymous):

\[\frac{ \cancel{6}3 }{ \cancel{2}(x+2)(x-1) }=\frac{ 3 }{ (x+2)(x-1) }\]

OpenStudy (anonymous):

the simplified expression...